OpenStudy (anonymous):

determine if a triangle DEF with coordinates D(2,1), E(3,5), and F(6,2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral ? be specific.

OpenStudy (anonymous):

will be given medal

OpenStudy (anonymous):

please someone help me!!

OpenStudy (anonymous):

whoever helps will be given a medal!!

OpenStudy (anonymous):

@zepdrix please help

zepdrix (zepdrix):

It's equilateral if all sides are the same length. So let's use the `distance formula` to figure out the length of each side. Calculating the distance from D to E,\[\Large\rm d_{1}=\sqrt{(3-2)^2+(5-1)^2}\]Understand how I plugged those values in?

OpenStudy (anonymous):

yes i do

zepdrix (zepdrix):

So what do you get for the length of side DE?

OpenStudy (anonymous):

sqrt 17

zepdrix (zepdrix):

\[\Large\rm \overline {DE}=\sqrt{17}\]Ok great! Understand how to check the lengths of the other sides? :o

OpenStudy (anonymous):

yes i get that but then how do i answer the second part of the question?

zepdrix (zepdrix):

Well you have to answer the first part before you can answer the second part. If they're all the same length, then you're done. If they're not, we can think about it. But figure out the lengths first! :o

OpenStudy (anonymous):

ok im doing that now :)

OpenStudy (anonymous):

ok so i got e to f sqrt 18 and d to f sqrt 17

OpenStudy (anonymous):

so its not equilateral

zepdrix (zepdrix):

\[\Large\rm \overline {DE}=\sqrt{17},\qquad \overline{EF}=\sqrt{18},\qquad \overline{FD}=\sqrt{17}\]Ok cool I got the same values. So it's currently Isosceles?

zepdrix (zepdrix):

So we need to fix this, hmmm

OpenStudy (anonymous):

yea the second part is where i was having a difficult time with

zepdrix (zepdrix):

This is the problem, When you `move a point`, it changes the length of `two sides`, not just one.

zepdrix (zepdrix):

So we don't want to try and turn the sqrt(18) into sqrt(17) by moving a point. We actually want to try and change the sqrt(17)'s into sqrt(18).

zepdrix (zepdrix):

One sec I gotta think about this one >.< hehe

OpenStudy (anonymous):

lol its fine no rush :)

zepdrix (zepdrix):

So this is what I'm thinking... We want to move our point E somewhere else to change our lengths of DE and FD. Let's call this new point (x,y). Again using the distance formula,\[\Large\rm \overline{DE}=\sqrt{(x-2)^2+(y-1)^2}\]And then set that equal to the length we want,\[\Large\rm \sqrt{18}=\sqrt{(x-2)^2+(y-1)^2}\]We can do the same with FD,\[\Large\rm \sqrt{18}=\sqrt{(x-6)^2+(y-2)^2}\]And that gives us a system of two equations. Mmmmmm that might not be the right approach, thinkinggggg >.<

OpenStudy (anonymous):

can we just change one coordinate to make it equal

zepdrix (zepdrix):

So the old coordinate was (3,5), you mean like leave the y at 5 and solve for the new x? I was trying to do that... I can't seem to make it work :( Grrr so confusingggg

zepdrix (zepdrix):

@jdoe0001 @dan815 @phi @campbell_st

OpenStudy (phi):

I would use the fact that an equilateral has 60 degree angles. we want to "rotate" radius EF to form a 60 degree angle at point E

OpenStudy (anonymous):

i dont understand..what do i do after

OpenStudy (anonymous):

@jdoe0001 can you try please

OpenStudy (phi):

one way to find where to move point E is to form the perpendicular bisector of side DF

OpenStudy (phi):

you can use algebra and find where that line intersects the circle. or (easier) we can use trigonometry to find the (x,y) values

OpenStudy (anonymous):

yes lets use trigonometry

OpenStudy (jdoe0001):

hmm

OpenStudy (phi):

the mid point of line segment DF is \[ (2,1) + \sqrt{17} (6-2,2-1) \\ (2,1) + \sqrt{17} (4,1) \] or, letting D be the origin for this exercise, \[ (4\sqrt{17},\sqrt{17}) \] the angle is inverse tangent of y/x = 1/4

OpenStudy (phi):

we get 14.036 degrees. Add 60º to get 74.036 now find the x and y coordinates (with D the origin) as sqr(17) cos(74.036) , sqr(17) sin(74.036) then add the coords of D to get the (x,y) values relative to the true origin.

OpenStudy (anonymous):

dont we find the measure of each angle and then shorten angle D?

OpenStudy (jdoe0001):

each angle is 60 degrees by defintion

OpenStudy (anonymous):

ok thank you guys so much

OpenStudy (phi):

There are different ways of doing this. Here is the idea I am using

OpenStudy (anonymous):

im doing this exam as the beginning of my geometry course and i dont think the answer should be that complex

OpenStudy (phi):

the idea is point G is point E moved a little bit (along the circle) so its length DG is sqr(17) we want to find the x and y offsets from pt G using R cos(angle) and R sin(angle) where R is sqr(17) and the angle is 60+ little angle

OpenStudy (phi):

If there is a short easy way to get a numerical answer, I am all ears. Perhaps they want a more "hand-waving" answer. e.g. rotate side DE so it forms a 60º angle with side DF

OpenStudy (jdoe0001):

maybe you're not required to get the (x,y) coordinates for F.... maybe you're just asked to give an explanation on what needs to be done to be an equilateral

OpenStudy (anonymous):

yea thats probably what they want

OpenStudy (anonymous):

well ill just put what @phi said thank you so much

OpenStudy (phi):

In geometry they teach you how to construct an equilateral triangle. Perhaps they want you to use this technique. See http://www.mathopenref.com/constequilateral.html